OpenStudy (anonymous):

Can you solve the following equation: abs(5x-3)=abs(3x+5)

OpenStudy (anonymous):

you'll have to solve these two equations...: 5x-3 = -1(3x+5) and 5x-3 = 3x+5

OpenStudy (anonymous):

what should I do next?

OpenStudy (anonymous):

According to your procedure, I got x=4. But the answer is not 4. Can you prove it?

OpenStudy (anonymous):

if you solve it correctly...

OpenStudy (cwrw238):

another way to do these is to square both sides and solve the quadratic

OpenStudy (anonymous):

i dont get is cwrw. can you write down the equations?

OpenStudy (anonymous):

 you actually have two solutions.... x=-1/4 and 4..